# appReferences

### v2

@@ Abstract
Learning from examples of success is an appealing approach to reinforcement learning but it presents a challenging exploration problem, especially for complex or long-horizon tasks. This work introduces value-penalized auxiliary control from examples (VPACE), an algorithm that significantly improves exploration in example-based control by adding examples of simple auxiliary tasks. For instance, a manipulation task may have auxiliary examples of an object being reached for, grasped, or lifted. We show that the naïve application of scheduled auxiliary control to example-based learning can lead to value overestimation and poor performance. We resolve the problem with an above-success-level value penalty. Across both simulated and real robotic environments, we show that our approach substantially improves learning efficiency for challenging tasks, while maintaining bounded value estimates. We compare with existing approaches to example-based learning, inverse reinforcement learning, and an exploration bonus. Preliminary results also suggest that VPACE may learn more efficiently than the more common approaches of using full trajectories or true sparse rewards. Videos, code, and datasets: https://papers.starslab.ca/vpace.
@@ I Introduction
Robotics presents a unique challenge to learning algorithms: because robotics data is costly to gather, ensuring that algorithms have high sample efficiency is crucial. In reinforcement learning and imitation learning, common approaches to improving sample efficiency incorporate a manually-defined dense reward function or demonstration trajectories of an expert completing the task, respectively. These additions, if they are available at all, can be highly biased or suboptimal. Sparse reward functions are less biased [1], but require significant exploration [2]. Sparse reward functions may also be unavailable, because they require accurate state estimation and the tuning of various parameters (e.g., how high must an object be lifted?).
We consider another form of feedback—example states of completed tasks, sometimes referred to as example-based control (EBC) [3]. Obtaining example states can be far less laborious than designing a reward function or gathering trajectories; practitioners can gather states from a distribution that represents a completed task without consideration of how the states are reached. However, just as in the case of sparse rewards, excluding information on how goal states are reached can lead to highly inefficient learning (e.g., an example of a loaded dishwasher provides no information about the long sequence of actions required to complete the task). We aim to address the following question: Can we improve the sample efficiency of example-based control?
@@ III-B Learning a Multitask Agent from Examples
We alleviate the challenging exploration problem of EBC by introducing auxiliary control from examples (ACE), an application of the scheduled auxiliary control framework [5, 6]. This hierarchical approach aims to improve exploration by learning from examples of simpler auxiliary tasks, in addition to a main task, to impose a curriculum where the completion of easier tasks provides guidance on how to complete harder ones [32]. Auxiliary tasks can be selected to have semantic meaning: in this work, the auxiliary tasks used are reach, grasp, lift, and release. The task definitions, and sometimes the examples themselves, can also be reusable between main tasks. Each task has a corresponding intention that learns from the given task-specific examples. ACE leverages a scheduler to sequentially choose and execute individual intention policies allowing for more diverse state coverage to facilitate faster learning of the main task. Crucially, because the method is off-policy, each intention can learn from data generated by any other intention, and all policy interaction data, regardless of which intention generated it, is stored in $\mathcal{B}$.
@@ IV-A2 Baselines
We consider many baselines, including an approach based on recursively classifying examples (RCE, [3]), a learned reward model for off-policy RL (DAC, [14]), a defined reward model for off-policy RL (SQIL, [15]), and a combination of the best performing baseline (SQIL) with a method that provides an exploration bonus to unseen data (SQIL+RND, [34, 35]). Finding that SQIL significantly outperformed the baselines, we use it as the default reward model for our main experiments, although VP and ACE can be added to any of the approaches described above. Specifically, in our results, VPACE refers to SQIL with both VP and ACE applied, while ACE refers to SQIL with ACE and without VP applied.
@@ IV-C3 Expert Data Quantity and $\lambda$ Value
Finally, to examine the robustness of our approach, we also compare various values for the quantity of examples of success (200 / task for main experiments) as well as the value of $\lambda$, which controls the strength of value penalization (10.0 for main experiments). Fig. 11 shows that changing the value of $\lambda$ to 1 and 100 has no effect on performance. Dropping from 200 to 100 examples has only a minor negative impact on performance in Insert, while the drop to 10 examples has a significant negative effect. However, even with only 10 examples, performance on Unstack-Stack still far exceeds all other baselines.

### v3

@@ Abstract
Common approaches to providing feedback in reinforcement learning are the use of hand-crafted rewards or full-trajectory expert demonstrations. Alternatively, one can use examples of completed tasks, but such an approach can be extremely sample inefficient. We introduce value-penalized auxiliary control from examples (VPACE), an algorithm that significantly improves exploration in example-based control by adding examples of simple auxiliary tasks and an above-success-level value penalty. Across both simulated and real robotic environments, we show that our approach substantially improves learning efficiency for challenging tasks, while maintaining bounded value estimates. Preliminary results also suggest that VPACE may learn more efficiently than the more common approaches of using full trajectories or true sparse rewards. Project site: https://papers.starslab.ca/vpace/.
@@ I Introduction
Example-based control (EBC) is a form of reinforcement learning in which example states of completed tasks are the only form of feedback [1]. Obtaining example states can be far less laborious than designing a reward function or gathering expert trajectories, and furthermore removes any possibility of using suboptimal expert trajectories. However, just as in the case of sparse rewards, excluding information on how goal states are reached can lead to highly inefficient learning (e.g., an example of a loaded dishwasher provides no information about the long sequence of actions required to complete the task). Can we improve the sample efficiency of example-based control?
@@ III-B Learning a Multitask Agent from Examples
We alleviate the challenging exploration problem of EBC by introducing auxiliary control from examples (ACE), an application of the scheduled auxiliary control framework [3, 4]. ACE leverages a scheduler to sequentially choose and execute individual intention policies allowing for more diverse state coverage to facilitate faster learning of the main task. Crucially, because the method is off-policy, each intention can learn from data generated by any other intention, and all policy interaction data, regardless of which intention generated it, is stored in $\mathcal{B}$.
@@ IV-A2 Baselines
We consider many baselines, including an approach based on recursively classifying examples (RCE, [1]), a learned reward model for off-policy RL (DAC, [12]), a defined reward model for off-policy RL (SQIL, [13]), and a combination of the best performing baseline (SQIL) with a method that provides an exploration bonus to unseen data (SQIL+RND, [31, 32]). Finding that SQIL significantly outperforms the baselines, we use it as the default reward model for our main experiments, although VP and ACE can be added to any of the approaches described above. Specifically, in our results, VPACE refers to SQIL with both VP and ACE applied, while ACE refers to SQIL with ACE and without VP applied.
@@ IV-C3 Expert Data Quantity and $\lambda$ Value
Finally, to examine the robustness of our approach, we also compare various values for the quantity of examples of success (200 / task for main experiments) as well as the value of $\lambda$, which controls the strength of value penalization (10.0 for main experiments). Fig. 11 shows that changing the value of $\lambda$ to 1 and 100 has no effect on performance, likely because it has no effect if the value constraint is not violated. Dropping from 200 to 100 examples has only a minor negative impact on performance in Insert, while the drop to 10 examples has a significant negative effect. However, even with only 10 examples, performance on Unstack-Stack still far exceeds all other baselines.
@@ V Conclusion
Acknowledgements
We gratefully acknowledge the Digital Research Alliance of Canada and NVIDIA Inc., who provided the GPUs used in this work through their Resources for Research Groups Program and their Hardware Grant Program, respectively. This work was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC). Jonathan Kelly gratefully acknowledges support from the Canada Research Chairs program.